[Brahms-l] Shift report 20050514 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat May 14 2005 - 07:49:08 EDT
Shifter leader was Michael
Shifters were 


Beam from Midnight-3am and then five until end of shift

No detector problems noticed

2.3 	3450 	B 	

runs 14620 14621  14622 14623 14624 14625  	
    Trig2 Trig3 Trig4 Trig5  
   548481 86582 57018 653227

May be missing events from 1 run since I forgot to press end run on shift log.


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat May 14 07:49:25 2005

This archive was generated by hypermail 2.1.8 : Sat May 14 2005 - 07:49:33 EDT